# Multiplicative polynomial equations in infinitely many variables

### v1

@@ Abstract.
This paper describes infinite sets of polynomial equations in infinitely many variables with the property that the existence of a solution or even an approximate solution for every finite subset of the equations implies the existence of a solution for the infinite set of equations.
@@
Systems of polynomial equations, systems of linear equations, approximate solutions, infinitely many variables, finite intersection property.
@@ 1. Finitely many implies infinite many
In mathematics there are theorems asserting that, for certain classes of equations, if every finite subset of an infinite set of the equations has a solution, then the infinite set of equations has a solution. For example, if every finite subset of an infinite set of linear equations in $n$ variables (that is, equations of the form $\sum_{j=1}^{n}a_{i,j}x_{j}=b_{i}$) has a solution, then the infinite set of linear equations in $n$ variables has a solution. For infinite sets of linear equations in infinitely many variables, in which each equation contains only finitely many variables (that is, equations of the form $\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}$ with $a_{i,j}\neq 0$ for only finitely many $j$), a “finitely solvable implies infinitely solvable” theorem is also true. For a survey of related results, see Nathanson [4].
@@
Let $(p,q)$ be a conjugate pair with $p>1$ and $q>1$ and let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$, Consider the linear equation in infinitely many variables
@@
Here is one solution. For all $r\in\mathbf{N}$, the sequence $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ defined by
@@
If $\mathbf{x}_{r}$ is any solution of the finite set of the first $r$ equations, then
$1=\sum_{j=r}^{\infty}a_{j}x_{j}=\sum_{j=1}^{\infty}a_{r,j}x_{j}=\|\mathbf{a}_{r}\mathbf{x}_{r}\|_{1}\leq\|\mathbf{a}_{r}\|_{p}\|\mathbf{x}_{r}\|_{q}$
@@
Let $(p,q)$ be a conjugate pair. Let $I$ be an infinite set and let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ for all $i\in I$. Consider the linear equation in infinitely many variables
@@
Let $M>0$. If, every finite subset $S$ of I and for every $\varepsilon>0$, the finite set of linear inequalities
@@
In this paper we generalize Theorem 2 to certain infinite sets of polynomial equations in infinitely many variables and prove (this is our main result) that the existence of norm-bounded or even sequentially bounded approximate solutions to all finite subsets of the set of polynomial equations is sufficient to guarantee the existence of an exact solution to the infinite set (Theorems 4 and 5). The theorems on linear equations are special cases of the polynomial results.
Our results apply to polynomials with coefficients and solutions in the field of real numbers and also in the field of complex numbers. However, the “finitely many implies infinitely many” paradigm is not true for all subfields of the complex numbers. It is not true, for example, for the field $\mathbf{Q}$ of rational numbers. Let $(b_{i})_{i\in I}$ be an infinite sequence of irrational numbers and consider the infinite set of linear equations $\{x_{i}=b_{i}:i\in I\}$. For every $\varepsilon>0$, every finite subset and, indeed, the infinite set of the equations has an approximate solution in $\mathbf{Q}$, but neither the infinite set nor any nonempty finite subset of the equations has an exact solution in $\mathbf{Q}$.
@@ 2. Polynomials in infinitely many variables
$\sum_{j=1}^{\infty}\left|a_{j}x_{j}^{d}\right|=\left\|\mathbf{a}\mathbf{x}^{d}\right\|_{1}\leq\left\|\mathbf{a}\right\|_{q/(q-d)}\left\|\mathbf{x}^{d}\right\|_{q/d}.$
@@
This is a finite set of cardinality $\binom{D}{k}$. The set $\bigcup_{k=1}^{D}\mathcal{D}_{k}$ has cardinality $\sum_{k=1}^{D}\binom{D}{k}=2^{D}-1$. For all $\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$ and $J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$, we define the monomial
(3) $x_{J}^{\Delta}=x_{j_{1}}^{d_{1}}x_{j_{2}}^{d_{2}}\cdots x_{j_{k}}^{d_{k}}$
@@
For example, let $D=3$. There are seven sets of monomials. For $k=1$ and $J=(j_{1})\in\mathbf{N}^{1}$, we have
@@
Monomials do not have a unique representation in the form $x_{J}^{\Delta}$. For example, with $D=3$, let $k=2$ and $J=(1,1)\in\mathbf{N}^{2}$. If $\Delta=(1,2)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}x_{1}^{2}=x_{1}^{3}$. If $\Delta=(2,1)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}^{2}x_{1}=x_{1}^{3}$. If $D=3$, $k=1$, $J=(1)\in\mathbf{N}^{1}$, and $\Delta=(3)\in\mathcal{D}_{1}$, then $x_{J}^{\Delta}=x_{1}^{3}$.
We do not need unique representation of monomials, but would have unique representation if we considered only $k$-tuples $(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$ such that $j_{1}<j_{2}<\cdots<j_{k}$.)
We consider only polynomials with zero constant term. In this paper, “polynomial” means polynomial with zero constant term.
A polynomial of degree at most $D$ in infinitely many variables $x_{1},x_{2},x_{3},\ldots$ is an infinite series of the form
@@
with coefficients $a_{\Delta,J}\in\mathbf{R}$. For example, a polynomial of degree at most 3 in infinitely many variables is of the form
@@ 3. Multiplicative polynomials
This is a finite sum, and so the series $P(\mathbf{x})$ converges absolutely for all $\mathbf{x}\in\ell^{q}$. Thus, for all $b\in\mathbf{R}$, the multiplicative polynomial equation
@@
is well-defined, and we can ask if this equation has a solution $\mathbf{x}\in\ell^{q}$.
Here is an example of a multiplicative polynomial. Let $1\leq d\leq D<q\leq\infty$ and so
$1\leq\frac{q}{q-1}\leq\frac{q}{q-d}\leq\frac{q}{q-D}.$
For all $s\in\mathbf{C}$ with $\Re(s)>(q-d)/q$, the Riemann zeta function
$\zeta\left(\frac{qs}{q-d}\right)=\sum_{j=1}^{\infty}\frac{1}{j^{qs/(q-d)}}=\sum_{j=1}^{\infty}\left(\frac{1}{j^{s}}\right)^{q/(q-d)}$
converges absolutely. Consider the sequence
$\mathbf{a}_{d}=\left(\frac{1}{j^{s}}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
and the polynomial
$(\mathbf{a}_{d},\mathbf{x}^{d})=\sum_{j=1}^{\infty}\frac{x_{j}^{d}}{j^{s}}.$
Let
$\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}\operatorname{\qquad\text{and}\qquad}J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}.$
For all $h\in\{1,\ldots,k\}$, let $s_{h}\in\mathbf{C}$ satisfy $\Re(s_{h})>(q-d)/q$ and let $\mathbf{a}_{d_{h}}=\left(j^{-s_{h}}\right)_{j=1}^{\infty}$. Then
The associated multiplicative polynomial of degree at most $D$ is
$P(\mathbf{x})$ $=P(\mathbf{x},s_{1},\ldots,s_{h})$
For $D=1$ and $q=2$ and for complex numbers $s$ with $\Re(s)>1/2$, the linear polynomial in infinitely many variables is the Dirichlet series
$P(\mathbf{x},s)=\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s}}.$
@@
For all $i\in I$, the sequences $\mathbf{a}_{i,d}$ determine the multiplicative polynomial equation
@@ 4. Approximate finite implies exact infinite
The set $\{P_{i}(\mathbf{x})=b_{i}:i\in S\}$ of polynomial equations in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ has an approximate solution if, for every $\varepsilon>0$, there exists a sequence $\mathbf{x}_{\varepsilon}$ such that $|P_{i}(\mathbf{x}_{\varepsilon})-b_{i}|\leq\varepsilon$ for all $i\in S$. We shall prove that an infinite set of multiplicative polynomial equations has an exact solution if every finite subset of the equations has a norm-bounded approximate solution. This is the main result in this paper and immediately implies Theorem 3.
@@
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. For all $k\in\mathbf{N}$, let
@@
. For all $i\in I$, the finite set of sequences $\{\mathbf{a}_{i,d}\}_{d=1}^{D}$ determines the multiplicative polynomial equation
@@
$=b_{i}.$
Let $M>0$. If, for every $\varepsilon>0$ and every finite subset $S$ of I, the finite set of polynomial inequalities
@@
The set
$X_{q,M}=\left\{\mathbf{x}\in\ell^{q}:\|\mathbf{x}\|_{q}\leq M\right\}$
is a compact subset of the topological space $\Omega$.
@@
Let $(p,q)$ be a conjugate pair and let $\mathbf{a}=(a_{j})_{j=1}^{\infty}\in\ell^{p}$. Let $M>0$. The linear functional $f$ on $X_{q,M}$ defined by
$f(\mathbf{x})=(\mathbf{a},\mathbf{x})=\sum_{j=1}^{\infty}a_{j}x_{j}$
@@
Let $U$ be an open subset of $\mathbf{R}$ or $\mathbf{C}$. We shall prove $f^{-1}(U)$ is open in $X_{q,M}$ as a subspace of $\Omega$.
If $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in f^{-1}(U)$, then $f(\mathbf{x})\in U$ and there exists $\varepsilon>0$ such that $U$ contains the open set
$\left\{t:|t-f(\mathbf{x})|<\varepsilon\right\}\subseteq U.$
@@
for all $N\geq N_{\mathbf{a}}$. Because $\mathbf{x}\in\ell^{q}$, the series $\sum_{j=1}^{\infty}|x_{j}|^{q}$ converges and there is an integer $N_{\mathbf{x}}$ such that
@@
The set
$V^{\prime}=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in\Omega:|y_{j}-x_{j}|<\delta\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
@@
=$=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in X_{q,M}:|y_{j}-x_{j}|<\delta\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
is an open neighborhood of $\mathbf{x}$ in $X_{q,M}$. We shall prove that $V\subseteq f^{-1}(U)$.
@@
$\tilde{y_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ y_{j}&\text{if $j\geq N+1$}\end{cases}$
@@
$\tilde{\mathbf{x}}=(\tilde{x_{j}})_{j=1}^{\infty}\in\ell^{q}\operatorname{\qquad\text{and}\qquad}\tilde{\mathbf{y}}=(\tilde{y_{j}})_{j=1}^{\infty}\in\ell^{q}.$
@@
We define
$\tilde{a_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ a_{j}&\text{if $j\geq N+1$}\end{cases}$
@@
$f(\mathbf{y})-f(\mathbf{x})$ $=(\mathbf{a},\mathbf{y}-\mathbf{x})=\sum_{j=1}^{\infty}a_{j}(y_{j}-x_{j})$
$=\sum_{j=1}^{N}a_{j}(y_{j}-x_{j})+\sum_{j=N+1}^{\infty}a_{j}(y_{j}-x_{j}).$
Applying the Hölder and Minkowski inequalities, we obtain
$\left|f(\mathbf{y})-f(\mathbf{x})\right|$ $\leq\sum_{j=1}^{N}|a_{j}|\ |y_{j}-x_{j}|+\sum_{j=N+1}^{\infty}|a_{j}(y_{j}-x_{j})|$
@@
Therefore, $f(\mathbf{y})\in U$ and $V\subseteq f^{-1}(U)$. Thus, the set $f^{-1}(U)$ is open in $X_{q,M}$ and the function $f(\mathbf{x})=(\mathbf{a},\mathbf{x})$ is continuous on $X_{q,M}$. This completes the proof. ∎
@@
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. By Lemma 2, the linear functionals $\left(\mathbf{a}_{i,d_{1}},\mathbf{x}^{d_{1}}\right)$ are continuous on $X_{q,M}$. Finite sums of finite products of continuous functions are continuous, and so the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{q,M}$ for all $i\in I$. It follows that, for all $i\in I$ and $\varepsilon>0$, the approximation set
@@
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. The set $X_{\mathbf{m}}$ is a closed subset of $X_{q,M}$ and so $X_{\mathbf{m}}$ is compact. By Lemma 2, for all $\mathbf{a}\in\ell^{p}$, the linear functional $f(\mathbf{x})=(\mathbf{a},\mathbf{x})$ is continuous on $X_{q,M}$, and so its restriction to $X_{\mathbf{m}}$ is continuous. It follows that the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{\mathbf{m}}$ for all $i\in I$, and so the approximation set

### v2

# Linear equations and multiplicative polynomial equations in infinitely many variables

@@ Abstract.
This paper describes infinite sets of linear and polynomial equations in infinitely many variables with the property that the existence of a solution or even an approximate solution for every finite subset of the equations implies the existence of a solution for the infinite set of equations.
@@
Systems of polynomial equations, systems of linear equations, approximate solutions, infinitely many variables, finite intersection property, Dirichlet series.
@@ 1. Finitely many implies infinite many
In mathematics there are theorems asserting that, for certain classes of equations, if every finite subset of an infinite set of the equations has a solution, then the infinite set of equations has a solution. For example, if every finite subset of an infinite set of linear equations in $n$ variables (that is, equations of the form $\sum_{j=1}^{n}a_{i,j}x_{j}=b_{i}$) has a solution, then the infinite set of linear equations in $n$ variables has a solution. For infinite sets of linear equations in infinitely many variables, in which each equation contains only finitely many variables (that is, equations of the form $\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}$ such that, for each $i$, we have $a_{i,j}\neq 0$ for only finitely many $j$), a “finitely solvable implies infinitely solvable” theorem is also true. For a survey of related results, see Nathanson [4].
@@
Let $(p,q)$ be a conjugate pair with $p>1$ and $q>1$ and let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$. Consider the linear equation in infinitely many variables
@@
Here is one solution of the triangular system. For all $r\in\mathbf{N}$, the sequence $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ defined by
@@
More generally, if $\mathbf{x}_{r}$ is any solution of the first $r$ equations of the triangular system, then
$1=\sum_{j=r}^{\infty}a_{j}x_{j}=\sum_{j=1}^{\infty}a_{r,j}x_{j}\leq\|\mathbf{a}_{r}\mathbf{x}_{r}\|_{1}\leq\|\mathbf{a}_{r}\|_{p}\|\mathbf{x}_{r}\|_{q}$
@@
Let $(p,q)$ be a conjugate pair. Let $I$ be an infinite set. For all $i\in I$, let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ and consider the linear equation in infinitely many variables
@@
Let $M>0$. If, for every finite subset $S$ of I and every $\varepsilon>0$, the finite set of linear inequalities
@@
In this paper we generalize Theorem 2 to certain infinite sets of polynomial equations in infinitely many variables and prove that the existence of norm-bounded or even sequentially bounded approximate solutions to all finite subsets of the set of polynomial equations is sufficient to guarantee the existence of an exact solution to the infinite set (Theorems 4 and 5). The theorem on linear equations is a special case of the polynomial results.
Our theorems apply to polynomials with coefficients and solutions in the field of real numbers and also in the field of complex numbers. However, the “finitely many implies infinitely many” paradigm is not true for all subfields of the complex numbers. For example, the paradigm fails for the field $\mathbf{Q}$ of rational numbers. Let $(b_{i})_{i\in I}$ be an infinite sequence of irrational numbers and consider the infinite set of linear equations $\{x_{i}=b_{i}:i\in I\}$. For every $\varepsilon>0$, every finite subset of the equations and, indeed, the infinite set of equations have approximate solutions in $\mathbf{Q}$, but neither the infinite set nor any nonempty finite subset of the equations has an exact solution in $\mathbf{Q}$.
@@ 2. Polynomials in infinitely many variables
$\sum_{j=1}^{\infty}\left|a_{j}x_{j}^{d}\right|=\left\|\mathbf{a}\mathbf{x}^{d}\right\|_{1}\leq\left\|\mathbf{a}\right\|_{q/(q-d)}\left\|\mathbf{x}^{d}\right\|_{q/d}<\infty.$
@@
This is a finite set of cardinality $\binom{D}{k}$. The set $\bigcup_{k=1}^{D}\mathcal{D}_{k}$ has cardinality $\sum_{k=1}^{D}\binom{D}{k}=2^{D}-1$. For all
$\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$
and
$J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$
we define the monomial
$x_{J}^{\Delta}=x_{j_{1}}^{d_{1}}x_{j_{2}}^{d_{2}}\cdots x_{j_{k}}^{d_{k}}$
@@
For example, let $D=3$ and $k\in\{1,2,3\}$. There are $17=2^{3}-$ 1 sets of monomials. For $k=1$ and $J=(j_{1})\in\mathbf{N}^{1}$, we have
@@
Monomials do not have a unique representation in the form $x_{J}^{\Delta}$. For example, with $D=3$, let $k=2$ and $J=(1,1)\in\mathbf{N}^{2}$. If $\Delta=(1,2)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}x_{1}^{2}=x_{1}^{3}$. If $\Delta=(2,1)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}^{2}x_{1}=x_{1}^{3}$. If $D=3$, $k=1$, $J=(1)\in\mathbf{N}^{1}$, and $\Delta=(3)\in\mathcal{D}_{1}$, then $x_{J}^{\Delta}=x_{1}^{3}$. Note that we do not need unique representation of monomials, but we would have unique representation if we considered only $k$-tuples $(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$ such that $j_{1}<j_{2}<\cdots<j_{k}$.)
We consider only polynomials with zero constant term. In this paper, “polynomial” means polynomial with zero constant term. By moving constants from one side of an equation to the other side, it is always possible to transform an equation that includes polynomials with nonzero constant terms to an equivalent equation in which no polynomial has a nonzero constant term.
A polynomial of degree at most $D$ in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ is a formal power series of the form
@@
with real or complex coefficients $a_{\Delta,J}$. For example, a polynomial of degree at most 3 in infinitely many variables is of the form
@@ 3. Multiplicative polynomials
This is a finite sum, and so the series $P(\mathbf{x})$ converges absolutely for all $\mathbf{x}\in\ell^{q}$.
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. A $q$-multiplicative polynomial of degree at most $D$ or, simply, a multiplicative polynomial, is a polynomial of degree at most $D$ in infinitely many variables of the form
$P(\mathbf{x})=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right)$
where
$\mathbf{a}_{d}=\left(a_{d,j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
for all $d\in\{1,\ldots,D\}$. The infinite series converges for all $\mathbf{x}\in\ell^{q}$.
A multiplicative polynomial equation is an equation of the form
@@
where $P(\mathbf{x})$ is a multiplicative polynomial. We ask if this equation has a solution $\mathbf{x}\in\ell^{q}$.
@@
For all $i\in I$, consider the multiplicative polynomial equation
@@ 4. Approximate finite implies exact infinite
The set $\{P_{i}(\mathbf{x})=b_{i}:i\in S\}$ of polynomial equations in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ has an approximate solution if, for every $\varepsilon>0$, there exists a sequence $\mathbf{x}_{\varepsilon}$ such that $|P_{i}(\mathbf{x}_{\varepsilon})-b_{i}|\leq\varepsilon$ for all $i\in S$. Theorem 4 states that an infinite set of multiplicative polynomial equations has an exact solution if every finite subset of the equations has a norm-bounded approximate solution. This is the main result of this paper and immediately implies Theorems 1, 2, and 3.
The theorem and the lemmas in this section are valid in both the real and complex fields.
@@
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. For all $k\in\{1,2,\ldots,D\}$, let
@@
For $\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$ and $J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$, define the multiplicative coefficient
@@
For all $i\in I$, the finite set of sequences $\{\mathbf{a}_{i,d}\}_{d=1}^{D}$ determines the multiplicative polynomial
@@
$=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{i,d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{i,d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{i,d_{k}},\mathbf{x}^{d_{k}}\right).$
Let $M>0$ and $\mathbf{b}=(b_{i})_{i\in I}$. If, for every $\varepsilon>0$ and every finite subset $S$ of $I$, the finite set of polynomial inequalities
@@
The set $X_{q,M}$ is a compact subset of the topological space $\Omega$.
@@
Let $d$ be a positive integer, $d<q\leq\infty$, and let $p=q/(q-d)$. Let $\mathbf{a}=(a_{j})_{j=1}^{\infty}\in\ell^{p}$. Let $M>0$. The function $f_{d}$ on $X_{q,M}$ defined by
$f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})=\sum_{j=1}^{\infty}a_{j}x_{j}^{d}$
@@
If $\mathbf{x}\in X_{q,M}$, then $\mathbf{x}^{d}\in\ell^{q/d}$. Because $(p,q/d)$ is a conjugate pair, the infinite series $\sum_{j=1}^{\infty}a_{j}x_{j}^{d}$ converges absolutely and the function $f_{d}(\mathbf{x})$ is well-defined.
Let $U$ be an open subset of $\mathbf{R}$ or $\mathbf{C}$. We shall prove $f_{d}^{-1}(U)$ is open in $X_{q,M}$ as a subspace of $\Omega$.
Let $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in f_{d}^{-1}(U)$. Because $f_{d}(\mathbf{x})\in U$, there exists $\varepsilon>0$ such that $U$ contains the open set
$\left\{t:|t-f_{d}(\mathbf{x})|<\varepsilon\right\}.$
@@
for all $N\geq N_{\mathbf{a}}$. Because $\mathbf{x}\in\ell^{q}$, the series $\sum_{j=1}^{\infty}\left|x_{j}\right|^{q}$ converges and there is an integer $N_{\mathbf{x}}$ such that
@@
Choose $\delta^{\prime}>0$ such that $|y_{j}-x_{j}|<\delta^{\prime}$ implies $|y_{j}^{d}-x_{j}^{d}|<\delta$ for all $j\in\{1,2,\ldots,N\}$. The set
$V^{\prime}=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in\Omega:|y_{j}-x_{j}|<\delta^{\prime}\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
@@
=$=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in X_{q,M}:|y_{j}-x_{j}|<\delta^{\prime}\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
is an open neighborhood of $\mathbf{x}$ in $X_{q,M}$. We shall prove that $V\subseteq f_{d}^{-1}(U)$.
@@
$\tilde{y_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ y_{j}&\text{if $j\geq N+1$.}\end{cases}$
Then
$\tilde{\mathbf{x}}=(\tilde{x_{j}})_{j=1}^{\infty}\in\ell^{q}\operatorname{\qquad\text{and}\qquad}\tilde{\mathbf{y}}=(\tilde{y_{j}})_{j=1}^{\infty}\in\ell^{q}$
@@
Thus, $\tilde{\mathbf{x}}\in X_{q,M}$ and $\tilde{\mathbf{y}}\in X_{q,M}$. We define
$\tilde{a_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ a_{j}&\text{if $j\geq N+1$.}\end{cases}$
@@
$\tilde{\mathbf{a}}=(\tilde{a_{j}})_{j=1}^{\infty}\in\ell^{p}$
@@
$f_{d}(\mathbf{y})-f_{d}(\mathbf{x})$ $=(\mathbf{a},\mathbf{y}^{d})-(\mathbf{a},\mathbf{x}^{d})$
$=\sum_{j=1}^{\infty}a_{j}y_{j}^{d}-\sum_{j=1}^{\infty}a_{j}x_{j}^{d}=\sum_{j=1}^{\infty}a_{j}\left(y_{j}^{d}-x_{j}^{d}\right)$
$=\sum_{j=1}^{N}a_{j}(y_{j}^{d}-x_{j}^{d})+\sum_{j=N+1}^{\infty}a_{j}(y_{j}^{d}-x_{j}^{d}).$
Applying the triangle, Hölder, and Minkowski inequalities, we obtain
$\left|f_{d}(\mathbf{y})-f_{d}(\mathbf{x})\right|$ $\leq\sum_{j=1}^{N}|a_{j}|\ |y_{j}^{d}-x_{j}^{d}|+\sum_{j=N+1}^{\infty}|a_{j}(y_{j}^{d}-x_{j}^{d})|$
@@
Therefore, $f_{d}(\mathbf{y})\in U$ and $V\subseteq f_{d}^{-1}(U)$. Thus, the set $f_{d}^{-1}(U)$ is open in $X_{q,M}$ and the function $f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})$ is continuous on $X_{q,M}$. This completes the proof. ∎
@@
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. By Lemma 2, the functions $\left(\mathbf{a}_{i,d_{i}},\mathbf{x}^{d_{i}}\right)$ are continuous on $X_{q,M}$ for all $d_{i}\in\{1,\ldots,D\}$. Finite sums of finite products of continuous functions are continuous, and so the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{q,M}$ for all $i\in I$. It follows that, for all $i\in I$ and $\varepsilon>0$, the approximation set
@@
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. By Lemma 2, for all $d\in\{1,\ldots,D\}$ and for all $\mathbf{a}\in\ell^{q/(q-d)}$, the function $f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})$ is continuous on $X_{q,M}$ and so its restriction to the compact set $X_{\mathbf{m}}$ is continuous. It follows that the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{\mathbf{m}}$ for all $i\in I$, and so the approximation set
@@
$\bigcap_{\begin{subarray}{c}i\in I\\ \varepsilon>0\end{subarray}}F_{i,\varepsilon}=\left\{\mathbf{x}\in X_{\mathbf{m}}:P_{i}(\mathbf{x})=b_{i}\text{ for all $i\in I$}\right\}$